### v1

@@ Abstract.
ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Applied computing Computer-assisted instruction
@@ 1. Introduction
In many low-resource schools across the Global South, teachers face severe constraints—teaching multiple subjects and grades in overcrowded classrooms with limited materials and digital tools. Lesson planning under these conditions is critical for shaping classroom structure, learning goals, and student outcomes (Farrell, 2002), yet remains a solitary, time-consuming task with minimal support. These structural challenges call for collaborative systems that support teacher work. Viewed through a CSCW lens, lesson planning is not just an individual task but a collaborative practice shaped by institutional mandates, peer networks, and now, algorithmic tools. It supports both instructional design and bureaucratic documentation—serving as written proof of teaching preparation and curricular compliance required by school administrators and education authorities (Farrell, 2002). Large Language Models (LLMs) introduce new possibilities by generating curriculum-aligned content across subjects and languages (Rashid et al., 2024a). Yet their use raises critical concerns: while they produce fluent text, they often fall short on factual accuracy, cultural sensitivity, and contextual alignment (Gallegos et al., 2024; Huang et al., 2024). Their effectiveness ultimately depends on how teachers interpret, adapt, and integrate these outputs into classroom practice (Bitkina et al., 2020b; OECD, 2024), especially in settings where technological stress compounds existing workloads (Varanasi et al., 2021).
@@
This paper examines the collaboration between teachers and AI in the context of lesson planning. We examine Shiksha copilot, a human-in-the-loop AI system deployed in Karnataka, India, between December 2024 and March 2025, designed to support grade 5–10 teachers create curriculum-aligned, contextual lesson plans (LPs) in English and Kannada. The system employs a structured workflow: LLMs generate initial content using Retrieval-Augmented Generation (RAG) frameworks; human curators then review these drafts for pedagogical quality, accuracy, and contextual relevance; teachers receive the curated materials and adapt them for classroom implementation. A subset of teachers also utilized supplementary AI features including curriculum-indexed chatbots for additional pedagogical support.
@@
Drawing on mixed-methods data from 1043 teachers—including system logs, interviews, and surveys—we analyze how teachers interact with AI-generated materials, how they adapt them to local classroom needs, and how this process reshapes their work. We approach this as a CSCW problem of human–AI collaboration in resource-constrained settings, where collaboration is not just about efficiency but about making teaching more contextually grounded, inclusive, and sustainable. To better understand how teachers collaborate with AI systems in everyday pedagogical work, we focus on three research questions:
@@
Our paper makes four contributions to CSCW: (1) We present the design and implementation of Shiksha copilot, a human-in-the-loop AI system that supports bilingual lesson planning in English and Kannada. By combining LLMs, RAG, and human curation, we offer insights into building educational AI systems for linguistically diverse contexts. (2) We provide empirical evidence of teacher–AI collaboration through analysis of system logs, interviews, and surveys. We document how teachers and curators evaluate, adapt, and contextualize AI-generated content, highlighting both the strengths and limits of current LLMs. (3) We show how AI tools reshape teacher workflows in Global South classrooms, addressing both instructional and administrative needs. Our findings trace how teachers integrate the tool, the challenges they face, and the organizational shifts that follow. (4) We advance CSCW theory on human–AI interaction by linking our findings to communities of practice and technology adoption in education. We also propose practical suggestions to address current LLM limitations in local-language content generation.
@@ 2. Related Work
Our work builds on two strands of CSCW and HCI research: (1) studies of teacher-facing technologies in the Global South that examine how digital systems are appropriated in resource-constrained educational settings, and (2) the emerging use of AI—especially LLMs—in teaching and pedagogy. These bodies of work foreground the sociotechnical nature of pedagogical work and provide foundation for understanding how emerging AI tools are integrated into everyday classroom practices and adapted by teachers in situated, often constrained environments.
@@ 2.1. Technology for Teachers in the Global South
Technology is often positioned as a pathway to educational equity, providing teachers with access to new instructional strategies and professional development opportunities (Purcell et al., 2013). However, such narratives frequently reflect techno-optimistic assumptions that overlook the need to align tools with local pedagogical practices, infrastructural constraints, and institutional realities (Mustafa et al., 2024; Chand et al., 2020; Warschauer and Ames, 2010). The One Laptop Per Child initiative, for instance, highlights how technological interventions can falter when they fail to consider teachers’ needs and classroom contexts (Warschauer and Ames, 2010; Wexler, 2021). Designing effective technological solutions, therefore, requires centering the experiences and constraints of teachers. Across India, sub-Saharan Africa, and Latin America, educators have shown a willingness to adopt digital tools and often adapt them creatively to local contexts, even in the face of infrastructural deficits, language barriers, and top-down implementation models that limit meaningful integration (Kaur and Singh, 2018; Nasreen and Chaudhary, 2018; Mustafa et al., 2024; Chand et al., 2020; Agyei, 2013). During the COVID-19 pandemic, Indian teachers used WhatsApp to share lessons and support peers - the Saharaline intervention was used as a structured support system, providing localized, ongoing pedagogical and emotional assistance through trained facilitators (Varanasi et al., 2024). In addition, Meghshala provided teachers with structured, multimedia LPs that they often adapted for linguistic and pedagogical fit (Reddy et al., 2022). Similarly, Open Educational Resources (OERs) have offered scalable ways to support instruction—particularly when paired with training or peer mentorship (Walker et al., 2023). In India, initiatives like TESS-India showed that localized, high-quality OERs can enhance teacher education and classroom practice (Perryman, 2013a). However, their success depends on adequate ICT skills, infrastructure, and culturally relevant content (Perryman, 2013a, b).
@@
Beyond tool access, teacher professional development remains a persistent challenge in the Global South. Although many teachers work in communities similar to their own, training programs often overlook the sociocultural and infrastructural complexities of classroom environments, including multilingual instruction, resource constraints, and large class sizes (Anwaruddin, 2016; Leach et al., 2005; Marcondes, 1999; Raina, 1999; Villegas-Reimers, 2003; Ramamurthy et al., 2020). Centralized curricula further restrict pedagogical autonomy, limiting teachers’ ability to adapt content to local needs (Ramamurthy et al., 2020; Finkelstein et al., 2013). Studies of large-scale initiatives in India have shown that professional development often fails to sustain meaningful pedagogical change, particularly when support is limited to tool onboarding or periodic mentoring (Ramamurthy et al., 2020; Varanasi et al., 2019). Prior work has emphasized that culturally responsive technology integration requires grounding design and training in teachers’ lived experiences, rather than assuming uniform uptake (Ogan et al., 2019; Uchidiuno et al., 2018). Our work builds on these insights by examining how teachers in Indian schools collaborate with an AI-based tool to create LPs. While prior interventions have explored how teachers adapt pre-authored content, the emergence of generative AI shifts the focus toward co-creating instructional materials in real time. In the following section, we examine recent work on AI tools for supporting teachers, with an emphasis on lesson planning, and the evolving dynamics of teacher–AI collaboration.
@@ 2.2. AI For Supporting Teachers
AI is increasingly being adopted in educational settings to assist teachers with a range of tasks, including grading, administrative support, generating instructional content, and providing personalized feedback (Wang et al., 2024; OECD, 2024; Dorottya Demszky, 2025). These applications promise to reduce teacher workload, improve instructional quality, and support differentiated instruction. However, their effectiveness depends heavily on curricular alignment, contextual integration, and teacher trust (Viberg et al., 2024; Bitkina et al., 2020a).
@@
Early tools like intelligent tutoring systems and automated rubric-based assessments laid the foundation for AI-assisted teaching (Anagnostopoulou et al., 2024). More recently, LLMs, such as ChatGPT, have emerged as generative tools capable of producing quizzes, comprehension exercises, and full LPs (Xiao et al., 2023; Baytak, 2024; Rashid et al., 2024b). LLM-generated LPs have been found to resemble those created by experienced educators, particularly benefiting novice teachers by scaffolding their planning and reducing preparation time (van den Berg and du Plessis, 2023; Karaman and Göksu, 2024). Yet, these benefits come with known limitations. LLMs often hallucinate facts, cite fabricated sources, and fail to adapt content to local pedagogical norms (Sridhar et al., 2023; Powell and Courchesne, 2024; Dornburg and Davin, 2024). Recent studies have identified four primary ways in which teachers seek support from LLM tools like ChatGPT (Keppler et al., 2025), and have also explored educator-centered perspectives on LLM-related harms and their wider implications (Harvey et al., 2025). Comparative evaluations reveal that AI-generated content tends to reinforce traditional teaching formats and rarely accommodates learner diversity or cultural specificity (Durmus, 2024; Clark and Van Kessel, 2024). In response, structured co-design frameworks have emerged. For instance, LessonPlanner embeds Gagné’s instructional principles into interactive UIs (Fan et al., 2024), while RAG frameworks refine LLM outputs through iterative feedback loops (Tan et al., 2025; Zheng et al., 2024; Mullins et al., 2024). Fine-tuned models and mega-prompting strategies that incorporate learner profiles, instructional goals, and teaching styles have also improved output fidelity (Karpouzis et al., 2024; Koenig, 2024). Despite these advances, most research and design efforts to use AI to improve teachers’ skills and capacity remain primarily concentrated in Western, English-dominant contexts (Bucchiarone et al., 2024). In multilingual and low-resource classrooms, these systems often struggle lack of localized datasets, insufficient language support, and unique challenges faced by the teachers (Sebastian et al., 2024).
@@
Studies of human–AI collaboration in educational contexts report limited impact (Küchemann et al., 2025), underscoring the need for more context-sensitive design that genuinely supports teachers’ practices and environments (Molenaar et al., 2022). In parallel, research emphasizes that effective AI integration requires not only improving AI content quality, but also fostering productive Teacher-AI Collaboration (TAC) (Sun et al., 2024). TAC is often framed through two perspectives: a replacement model where AI substitutes the teacher in delivering instruction or assessments, and an augmentation model that supports teachers with content creation, feedback, and personalization (Jeon et al., 2023; Baker and Inventado, 2016). Recent work has largely converged on the augmentation model, emphasizing that teachers remain irreplaceable due to their pedagogical insight and contextual expertise (Lawrence et al., 2024; Park and Park, 2024). These findings suggest that collaboration is most effective when teachers exercise design judgment and bring domain knowledge to bear. Teacher agency—the capacity to evaluate, adapt, or reject AI suggestions—is thus key to fostering meaningful collaboration (Jeon and Kim, 2022; Nazaretsky et al., 2022). TAC in the Global South often depends on a broader ecosystem involving designers, content curators, and institutional intermediaries. For example, Choi et al. (Choi et al., 2024) found that while an AI chatbot improved teacher confidence in Sierra Leone, infrastructural limitations and trust gaps hindered sustained use. In India, educators valued the time-saving potential of generative tools but faced usability barriers including long response times, absence of vernacular content, and misalignment with school routines (Sebastian et al., 2024; Santos et al., 2024). Our work extends this emerging literature by examining teacher-AI collaboration in Indian classrooms, where educators must adapt LLM outputs to multilingual contexts, assess pedagogical fit, and mediate their role as instructional decision-makers. In addition to exploring the questions of LLM-system accuracy, we explore how teachers embed AI into their everyday workflows and discuss the social, cultural, organizational, and infrastructural dimensions of teacher–AI interaction.
@@ 3. System Design and Implementation
Shiksha copilot was developed to support teachers in low-resource Indian classrooms by enhancing lesson planning workflows through an LLM-based system with human-in-the-loop curation and bilingual support. The system was developed in close collaboration with an implementation partner that runs several initiatives to strengthen public education in Karnataka, India. To inform the design, the partner conducted workshops with teachers to identify key pain points and opportunities where AI-based tools could offer meaningful support. These consultations surfaced four critical needs: alignment with the state-mandated curriculum, scaffolding for low-resource classrooms with limited digital infrastructure and learning materials, compliance with the 5E pedagogical format (Bybee et al., 2006; Duran and Duran, 2004) mandated by the Karnataka education department, and contextual adaptability for diverse teaching environments. In response to these inputs, the system was designed to prioritize curriculum-aligned content generation and support contextualized teaching across both English and Kannada mediums. Its design was grounded in the following principles:

### v2

@@ Abstract.
$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Applied computing Computer-assisted instruction
@@ 1. Introduction
In many low-resource schools across the Global South, teachers often work in environments with overcrowded classrooms, staff shortages, and heavy administrative duties. These pressures contribute to time scarcity, excessive workloads, and teacher burnout (Creagh et al., 2025; Parihar and Mahmood, 2016; Kim, 2019). These challenges are compounded when teachers are required to teach multiple subjects across different grade levels, with limited institutional and pedagogical support. In such contexts, developing lesson plans is a critical tool for structuring instruction and ensuring students meet learning goals. Yet, the very constraints teachers face make it especially difficult to create high-quality lesson plans (Yuan and Zhang, 2016; Dorji and Wangchuk, 2022), despite their proven role in improving teaching effectiveness and student outcomes (Piper et al., 2018). Beyond guiding instruction, lesson plans also serve a bureaucratic function, providing written evidence of preparation and curricular compliance required by school administrators and education authorities (Farrell, 2002). These structural constraints, combined with the critical role of lesson plans in both pedagogy and accountability, underscore the need for robust support systems to help teachers design and manage high-quality lessons.
Large Language Models (LLMs) have shown great promise for generating high-quality lesson plans across subjects and languages (Rashid et al., 2024a). LLM-based tools like Khanmigo [^1] https://www.khanmigo.ai/teachers and MagicSchool AI [^2] https://www.magicschool.ai/ have already made these capabilities available to teachers, providing them pedagogical support and ready-made templates to ease their workload. However, despite the promise and reach of these tools, significant knowledge gaps remain which are central to CSCW research. First, while LLMs produce fluent text, they often fall short on factual accuracy, cultural sensitivity, and contextual alignment (Gallegos et al., 2024; Huang et al., 2024), particularly for languages and contexts in the Global South (Gallegos et al., 2024; Mirza et al., 2024). Even when outputs are accurate, teachers must still interpret, adapt, and integrate these outputs into classroom practices (Bitkina et al., 2020b; OECD, 2024), a process that is shaped by institutional mandates and peer practices. Furthermore, we know little about how they perform in practice, particularly in Global South, where infrastructure, language, and pedagogical support differ sharply from Western contexts.
@@
This paper presents the design and development of an LLM-based lesson planning tool and examines how teachers in India used it in practice, focusing on collaboration, implementation challenges, and its influence on their work practices. We designed, deployed, and evaluated Shiksha Copilot, a human-in-the-loop AI system to support teachers create curriculum-aligned, contextual lesson plans (LPs) in English and their local language. The tool was deployed in Karnataka, India, between December 2024 and March 2025, to support grade 5–10 teachers to create LPs in English and Kannada. The system employs a structured workflow: LLMs generate initial content using Retrieval-Augmented Generation (RAG) frameworks; human curators then review these drafts for pedagogical quality, accuracy, and contextual relevance; teachers receive the curated materials and adapt them for classroom implementation. Teachers can also utilize supplementary AI features including curriculum-indexed chatbots for additional pedagogical support.
@@
Drawing on mixed-methods data from 1043 teachers—including system logs, interviews, and surveys—we analyze how teachers interact with AI-generated materials, how they adapt them to local classroom needs, and how this process reshapes their work. In particular, we ask:
@@
Building on our findings, we discuss what makes teacher–AI collaboration effective, emphasizing teachers’ active role as evaluators and adopters of AI-generated content. We highlight how AI-driven educational tools can be embedded within existing Communities of Practice to foster collective knowledge creation and ensure cultural and linguistic relevance. Finally, we reflect on the limitations of current LLMs for “low-resource” languages and argue for human-in-the-loop and community-driven approaches to improve quality, equity, and sustainability of AI in education. In sum, our paper makes three contributions to CSCW:
  • We present the design and implementation of Shiksha Copilot, a human-in-the-loop AI system that supports bilingual lesson planning, offering insights into building Human-AI systems for linguistically diverse educational contexts.
  • We provide empirical evidence of teacher–AI collaboration, demonstrating how teachers evaluate, adapt, and contextualize AI-generated content, highlighting both the strengths and limits of current LLMs.
  • We show how AI tools transform teacher workflows in Global South classrooms, influencing both instructional practices and administrative processes, and driving organizational change.
@@ 2. Related Work
Our work builds on two strands of CSCW and HCI research: (1) studies of teacher-facing technologies in resource-constrained educational settings in the Global South, and (2) the emerging use of AI, especially LLMs, in teaching and pedagogy. Together, these bodies of work provide a foundation for understanding how emerging AI tools are integrated into everyday classroom practices and adapted by teachers working in situated, often highly constrained, environments.
@@ 2.1. Technology for Teachers in the Global South
Technology is often positioned as a pathway to educational equity, providing teachers with access to new instructional strategies and professional development opportunities (Purcell et al., 2013). However, such narratives frequently reflect techno-optimistic assumptions that overlook the need to align tools with local pedagogical practices, infrastructural constraints, and institutional realities (Mustafa et al., 2024; Chand et al., 2020; Warschauer and Ames, 2010). The One Laptop Per Child initiative, for instance, highlights how technological interventions can falter when they fail to consider teachers’ needs and classroom contexts (Warschauer and Ames, 2010; Wexler, 2021). Designing effective technological solutions, therefore, requires centering the experiences and constraints of teachers.
Across India, sub-Saharan Africa, and Latin America, educators have shown a willingness to adopt digital tools and often adapt them creatively to local contexts, even in the face of infrastructural deficits, language barriers, and top-down implementation models that limit meaningful integration (Kaur and Singh, 2018; Nasreen and Chaudhary, 2018; Mustafa et al., 2024; Chand et al., 2020; Agyei, 2013). For example, during the COVID-19 pandemic, Indian teachers used WhatsApp-based helpline to receive pedagogical and emotional support through trained facilitators (Varanasi et al., 2024). Meghshala, a teacher-support organization, provided teachers with structured, multimedia-driven LPs that teachers could adapt to their classroom needs (Reddy et al., 2022). Similarly, Open Educational Resources (OERs) offered scalable ways to support instruction—particularly when paired with training or peer mentorship (Walker et al., 2023). In India, initiatives like TESS-India showed that localized, high-quality OERs can enhance teacher education and classroom practice (Perryman, 2013a). However, the success of these initiatives ultimately hinges on more than just content availability; it requires adequate infrastructure, teachers’ digital skills, and materials that are culturally and linguistically relevant to their classrooms (Perryman, 2013a, b).
@@
Beyond tool access, teacher professional development remains a persistent challenge in the Global South. Although many teachers work in communities similar to their own, training programs often overlook the sociocultural and infrastructural complexities of classroom environments, including multilingual instruction, resource constraints, and large class sizes (Anwaruddin, 2016; Leach et al., 2005; Marcondes, 1999; Raina, 1999; Villegas-Reimers, 2003; Ramamurthy et al., 2020). Centralized curricula further restrict pedagogical autonomy, limiting teachers’ ability to adapt content to local needs (Ramamurthy et al., 2020; Finkelstein et al., 2013). Prior work has emphasized that culturally responsive technology integration requires grounding design and training in teachers’ lived experiences, rather than assuming uniform uptake (Ogan et al., 2019; Uchidiuno et al., 2018). Our work builds on these insights by examining how teachers in Indian schools collaborate with an AI-based tool to create LPs. While prior interventions have explored how teachers adapt pre-authored content, the emergence of generative AI shifts the focus toward co-creating instructional materials in real time. In the following section, we examine recent work on AI tools for supporting teachers, with an emphasis on lesson planning, and the evolving dynamics of teacher–AI collaboration.
@@ 2.2. AI For Supporting Teachers
AI is increasingly being adopted in educational settings to assist teachers with a range of tasks, including grading, administrative support, generating instructional content, and providing personalized feedback (Wang et al., 2024; OECD, 2024; Dorottya Demszky, 2025). These applications promise to reduce teacher workload, improve instructional quality, and support differentiated instruction. However, their effectiveness depends heavily on curricular alignment, contextual integration, and teacher trust (Viberg et al., 2024; Bitkina et al., 2020a). At the same time, Salehi et al. (2024) caution that introducing AI into knowledge work can reconfigure labor and deskill practitioners. This is particularly relevant for teachers who make extensive use of AI in their knowledge work.
Early tools such as intelligent tutoring systems and automated rubric-based assessments laid the foundation for AI-assisted teaching (Anagnostopoulou et al., 2024). More recently, LLMs, such as ChatGPT, have emerged as generative tools capable of producing quizzes, comprehension exercises, and even complete LPs (Xiao et al., 2023; Baytak, 2024; Rashid et al., 2024b). Studies show that LLM-generated LPs often resemble those created by experienced educators and can be particularly helpful for novice teachers, offering pedagogical support and reducing preparation time (van den Berg and du Plessis, 2023; Karaman and Göksu, 2024).
Yet these benefits are accompanied by well-documented limitations. Research has shown that LLMs can hallucinate, generate fabricated citations, and struggle to adapt content to local pedagogical norms (Sridhar et al., 2023; Powell and Courchesne, 2024; Dornburg and Davin, 2024). Harvey et al. (2025), for example, examine educator-centered perspectives on LLM-related harms and their broader implications. Studies of human–AI collaboration in educational contexts report mixed or limited impact (Küchemann et al., 2025), highlighting the need for more context-sensitive design that genuinely supports teachers’ practices and environments (Molenaar et al., 2022). Comparative evaluations also show that AI-generated content often reinforces rigid teaching formats and rarely accommodates learner diversity or cultural specificity (Durmus, 2024; Clark and Van Kessel, 2024).
@@
In response to these challenges, researchers have developed structured co-design frameworks to better align AI outputs with classroom needs. For instance, LessonPlanner embeds Gagné’s instructional principles into interactive UIs (Fan et al., 2024), while RAG frameworks refine LLM outputs through iterative feedback loops (Tan et al., 2025; Zheng et al., 2024; Mullins et al., 2024). Fine-tuned models and mega-prompting strategies that incorporate learner profiles, instructional goals, and teaching styles have also improved output fidelity (Karpouzis et al., 2024; Koenig, 2024). However, despite these advances, most research and design efforts to use AI to improve teachers’ skills and capacity remain primarily concentrated in Western, English-dominant contexts (Bucchiarone et al., 2024). In multilingual and low-resource classrooms, these systems often struggle because of lack of localized datasets, insufficient language support, and unique challenges faced by the teachers (Sebastian et al., 2024).
Research also emphasizes that effective AI integration requires not only improving AI content quality, but also fostering productive Teacher-AI Collaboration (TAC) (Sun et al., 2024). TAC is often framed through two perspectives: a replacement model where AI substitutes the teacher in delivering instruction or assessments, and an augmentation model that supports teachers with content creation, feedback, and personalization (Jeon et al., 2023; Baker and Inventado, 2016). Building on this augmentation view, recent theory argues that how we frame AI—as a tool versus a collaborator—shapes whose agency is amplified. Satyanarayan and Jones (2024) propose treating humans and GenAI as co-agents whose compositional and interpretive roles can be dynamically reconfigured, with “design as delegation of constrained agency” and “ethics as care” guiding longer-term stewardship. For example, Shiksha Copilot, the teacher-AI collaborative system that we study, follows this co-agency stance, keeping teachers accountable decision-makers while letting the model assume bounded, negotiable roles in drafting and refinement.
@@
Empirical studies in the Global South further illustrate how TAC depends on broader ecosystems involving designers, content curators, and institutional intermediaries. For example, Choi et al. (Choi et al., 2024) found that while an AI chatbot improved teacher confidence in Sierra Leone, infrastructural limitations and trust gaps hindered sustained use. In India, educators valued the time-saving potential of generative tools but faced usability barriers including long response times, absence of vernacular content, and misalignment with school routines (Sebastian et al., 2024; Santos et al., 2024). Our work extends this emerging literature by examining teacher-AI collaboration in Indian classrooms, where educators must adapt LLM outputs to multilingual contexts, assess pedagogical fit, and mediate their role as instructional decision-makers. In addition to exploring the questions of LLM-system accuracy, we explore how teachers embed AI into their everyday workflows and discuss the social, cultural, organizational, and infrastructural dimensions of teacher–AI interaction.
@@ 3. System Design and Implementation
Shiksha Copilot was developed to support teachers in low-resource Indian classrooms by enhancing lesson planning workflows through an LLM-based system with human-in-the-loop curation and bilingual support. The system was developed in close collaboration with an implementation partner that runs several initiatives to strengthen public education in Karnataka, India for over 22 years. Guru was piloted in Karnataka as part of the organization’s larger vision of making quality teaching and learning resources accessible to English- and Kannada-medium public school teachers. Aside from Guru, the interventions designed by the partner organization are implemented on the ground by mentors, who serve as conduits between the organization’s program team and public school teachers. These mentors train the teachers and provide ongoing support across all programs. The same approach was followed for Guru. The costs for lesson plan curation, tool deployment with teachers, and LLM API usage were funded by the partner organization through philanthropic grants.
To inform the design, the partner conducted focus-group discussions with 29 teachers in batches to identify key pain points and opportunities where AI-based tools could offer meaningful support. The partner organization surfaced three critical needs through these consultations: alignment with the state-mandated curriculum, scaffolding for low-resource classrooms with limited digital infrastructure and learning materials, compliance with the 5E Instructional Model.
The 5E Instructional Model is a constructivist framework for teaching that guides learners through five phases: Engage, Explore, Explain, Elaborate, and Evaluate. In the Engage phase, teachers spark curiosity and connect prior knowledge to new ideas; in Explore, students investigate phenomena through inquiry and hands-on activities; in Explain, they articulate understanding while teachers formalize concepts; during Elaborate, students apply knowledge to new contexts; and in Evaluate, both teachers and students assess learning and reflect on progress (Polanin et al., 2024).
The 5E format is mandated by the Karnataka education department, and contextual adaptability for diverse teaching environments. In response to these inputs, the system was designed to prioritize curriculum-aligned content generation and support contextualized teaching across both English and Kannada mediums. Its design was grounded in the following principles: